A vertical cylinder tank of straight section area 'A1' is open to the air at its top and contains water up to depth 'h0'. A hole of area 'A2' is drilled in the bottom (A1> A2).
(A) Show in detail how to get to the formula below with the data provided.

h (t) = ((√h0) - (A2 / A1). (√g / 2) .t) ²

(B) How long does it take for the tank to empty after the hole is opened? (The answer must be based on A1, A2, h0 and g)
1
Expert's answer
2020-04-22T09:46:38-0400

a)


"dV=A_2vdt=A_2\sqrt{2gh}dt"

"dV=A_1dh"

"A_2\sqrt{2gh}dt=A_1dh"

"\sqrt{h_0}-\sqrt{h}=\frac{A_2}{A_1}\sqrt{\frac{g}{2}}t"

"h=\left(\sqrt{h_0}-\frac{A_2}{A_1}\sqrt{\frac{g}{2}}t\right)^2"

b)


"h=0\to \sqrt{h_0}=\frac{A_2}{A_1}\sqrt{\frac{g}{2}}t"

"t=\sqrt{\frac{2h_0}{g}}\frac{A_1}{A_2}"